What are the real and imaginary parts of CosZ?



1
Expert's answer
2022-02-08T11:20:21-0500

"\cos Z = \cos (x+iy) = \cos x \cos iy – \sin x \sin iy\\
= \cos x \cos h y – i \sin x \sin hy\\"

Here real part of "\cos Z" is "cos x \cos h y" and imaginary part of "\cos Z"

is "\sin x \sin hy\\" .